# Detection of Lensed Gravitational Waves from dark matter halos with deep learning

### v1

@@ 5 Deep Learning Architecture and Training Strategy
To evaluate the benefit of multiband observations, two network configurations were implemented. The first is a single branch network processing input waveforms from either DECIGO or ET individually. The second is a dual branch fusion network designed for multiple detector inputs, where each branch independently processes data from one detector using identical ResNet substructures. Deep feature maps from both branches are concatenated in feature space and fused via a fully connected layer to form a unified representation. This late fusion approach preserves detector specific feature learning while leveraging the complementarity between the low frequency (DECIGO) and mid to high frequency (ET) bands, thereby enhancing discrimination of both interference-diffraction patterns and geometric magnification effects (Fig. 6).
@@
6 Results and Analysis
This section evaluates the model’s classification performance under different detector setups and SNR conditions. The analysis is divided into two parts. First, using datasets that span the full SNR range of unlensed signals (0-$100$), we evaluate and compare the classification results of models trained on individual detectors (ET and DECIGO) with those trained on combined data from both detectors. This comparison helps to quantify the effect of using multiple detectors on overall performance. Second, we examine how classification accuracy varies with SNR within the multiple detector configuration, to better understand how signal strength influences detection sensitivity and the reliability of lens classification.
6.1 Single and Multiple Detector Performance
Using a comprehensive dataset spanning the 0-$100$ SNR range, we evaluated the model’s classification performance across five classes: PN, UL signals, and three representative lensing models (SIS, CIS, and NFW). Three input configurations were tested: single detector ET, single detector DECIGO, and the multiple detectors (ET+DECIGO). The dataset was divided into training and testing subsets in a 9:19{:}1 ratio. Each class contained $4\times 10^{4}$ samples, corresponding to $3.6\times 10^{4}$ training and $4\times 10^{3}$ testing samples per class, giving a total of $1.8\times 10^{5}$ training and $2\times 10^{4}$ testing samples.
@@ 6.1 Single and Multiple Detector Performance
$\mathrm{Micro\penalty 10000\ average}(M)=\frac{1}{C}\sum_{i=1}^{C}M_{i},$ (43)
@@
Comprehensive comparisons under different input configurations are summarised in Tables 3-5. The accuracy, average $F_{1}$, and average AUC trends are shown in Fig. 7, and the corresponding ROC curves and confusion matrices are presented in Fig. 8 and Fig. 9, respectively.
@@
The confusion matrix (Fig. 9(c)) exhibits strong diagonal dominance, with mutual misclassification between SIS and CIS below $2\%$, and NFW misclassification below $1\%$. The ROC curves (Fig. 8) show that the joint model completely encompasses both single detector models, with a micro average AUC approaching unity and individual class AUCs improving by about 0.100.10-$0.25$. This demonstrates a clear enhancement in both sensitivity and stability.
@@
Overall, based on the metrics in Tables 3-5 and the visual comparisons in Figs. 8-9, the ET+DECIGO joint model outperforms both single detector counterparts in accuracy, confusion structure, and robustness, confirming the physical and statistical advantages of multiband data fusion for detecting lensed GWs in noisy environments.
6.2 Performance of the Multiple Detector at Different SNR
To evaluate the robustness and sensitivity of the multiple detector (ET+DECIGO) model under varying SNR, the model was trained and evaluated separately within five SNR intervals: 55-2020, 2020-$40$, 4040-$60$, 6060-$80$, and 8080-$100$. Each interval included five classes (PN, UL, SIS, CIS, NFW) with a total of $2\times 10^{5}$ samples ($4\times 10^{4}$ per class). The data were divided into training and testing subsets with a 9:19{:}1 split. Model performance was quantified using Accuracy, Precision, Recall, $F_{1}$ score, and Macro Average metrics. Additionally, confusion matrices (Fig. 12) and ROC curves (Fig. 11) were plotted to visualise class separation across different SNR regimes. Quantitative metrics for each interval are summarised in Tables 6-10, and the evolution of accuracy, average $F_{1}$, and average AUC with increasing SNR is shown in Fig. 10.
@@ 6.2 Performance of the Multiple Detector at Different SNR
At the lowest SNR interval (55-2020), model performance is strongly degraded by noise contamination. The accuracy reaches only $0.6339$, and the macro averaged $F_{1}$ score is $0.6059$ (Table 6). While the “PN” class is perfectly identified ($F_{1}=1.000$), the “UL” and three lensed classes exhibit substantial confusion. The confusion matrix (Fig. 12(a)) shows that SIS and CIS suffer the most from mutual misclassification, with correct recognition rates of about $39\%$ and $15\%$, respectively, and a fraction of samples incorrectly labelled as NFW. This indicates that at low SNR, waveform interference features are heavily masked by noise, hindering the network’s ability to extract stable phase dependent structures. The ROC curves (Figs. 11(d), 11(e), and 11(f)) show that the AUC values for the three lensed classes remain in the range 0.840.84-$0.85$, suggesting partial robustness to noise but limited discriminative confidence. In this regime, the joint model’s advantage is not yet fully realised, being primarily constrained by the low SNR boundary of the input data.
@@
As the SNR increases to moderate levels (2020-$40$ and 4040-$60$), the model performance improves markedly. In the 2020-$40$ range, the accuracy rises to $0.7074$ (Table 7), and the macro averaged $F_{1}$ score reaches $0.6427$. Both SIS and NFW exhibit substantial gains relative to the low SNR regime, with $F_{1,\text{SIS}}$ improving from $0.4431$ to $0.5891$ and $F_{1,\text{NFW}}$ from $0.5496$ to $0.6620$. The CIS class, however, remains poorly classified, with $F_{1}=0.0215$.
At SNRs of 4040-$60$, performance increases sharply: accuracy reaches $0.9642$ (Table 8), and the macro averaged $F_{1}$ rises to $0.9634$. The five classes become nearly perfectly separated, with “PN” and “UL” achieving $F_{1}=1.000$. Among the lensed signals, $F_{1,\text{SIS}}=0.9170$, $F_{1,\text{CIS}}=0.9009$, and $F_{1,\text{NFW}}=0.9995$. The confusion matrix (Fig. 12(c)) exhibits clear diagonal dominance, with residual misclassifications below $20\%$. The ROC curves (Fig. 11) show AUC values exceeding $0.9$ for all classes, with an average AUC near $0.99$, confirming the model’s capacity to distinguish diffraction and geometric magnification features at moderate SNRs. The accuracy, average $F_{1}$, and average AUC in Fig. 10 displays a distinct transition around an SNR of $40$, indicating that multiple detector data fusion becomes particularly effective beyond this threshold.
@@
At high SNR (6060-$100$), model performance approaches saturation. In the 6060-$80$ SNR interval, accuracy reaches $0.9983$ (Table 9), and the macro averaged $F_{1}$ is $0.9982$, with both Precision and Recall for all classes nearly unity. The confusion matrix (Fig. 12(d)) is almost perfectly diagonal, showing only negligible misclassifications.
In the 8080-$100$ SNR range, accuracy slightly decreases to $0.9903$ (Table 10), mainly due to minor boundary overlaps among classes, but remains close to ideal. The $F_{1}$ scores are $F_{\text{SIS}}=0.9750$, $F_{1,\text{CIS}}=0.9767$, and $F_{1,\text{NFW}}=0.9994$. The ROC curves (Figs. 11(a) and 11(b)) show $\mathrm{AUC}\approx 1.000$, confirming that the model achieves near perfect discrimination at high SNR. The confusion matrix (Fig. 12(e)) further verifies precise classification across all five classes, with virtually no off diagonal errors.
The overall trend (Fig. 10) reveals a rapid convergence of model performance with increasing SNR—from $63.4\%$ accuracy in the 55-2020 range to $99.8\%$ in the 6060-$80$ interval—followed by stability at higher SNRs. These results confirm that the multiple detector configuration substantially enhances the extraction of interference and amplitude features at low to moderate SNRs, while maintaining calibrated probability outputs and precise classification at high SNR. Combining the results in Tables 6-10 and Figs. 11-12, it is evident that the ET+DECIGO model achieves consistently high accuracy and robustness across all SNR regimes, highlighting the crucial role of multiband information fusion in detecting lensed GWs under noise conditions.
7 Conclusion and Outlook
@@
Future research can further extend and refine this framework in several directions. In data generation, incorporating more diverse lens models, such as ellipsoidal potentials (hawken2009gravitational) and double lens systems (collett2016compound; vegetti2024strong), may enhance the adaptability and generalization of the model in complex lensing environments. On the algorithmic side, future studies could investigate network architectures that integrate time and frequency domain representations, or introduce attention mechanisms to better capture long timescale features (chen2023long; brauwers2021general). Joint task learning strategies may also enable simultaneous lens classification and parameter estimation, for example, predicting lens mass or impact parameter. The proposed approach can be further evaluated under more realistic observational conditions, including multiple overlapping sources, non stationary noise, and anisotropic detector responses, to examine its robustness and scalability. As third generation ground based detectors (ET) and space detectors (DECIGO, LISA (amaro2017laser), Taiji (ruan2020taiji), TianQin (luo2016tianqin)) continue to advance, multiband and multiband observations are expected to play an increasingly important role in gravitational wave astronomy. Within this context, the present framework offers a useful foundation for rapid identification of lensed GWs, inference of lens parameters, and statistical investigations related to dark matter and cosmology, contributing to the ongoing development of precision gravitational wave astrophysics.
@@ Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203004), the Fundamental Research Funds for the Central Universities Project (Grant No. 2024IAIS-ZD009), the National Natural Science Foundation of China (Grant Nos. 12575072 and 12347101), and the Natural Science Foundation of Chongqing (Grant No. CSTB2023NSCQ-MSX0103). This work made use of TensorFlow (Abadi_TensorFlow_Large-scale_machine_2015), PyCBC (alex_nitz_2024_10473621), GLoW_public (GLoW_public), and other open source Python packages.

### v2

@@ V Deep Learning Architecture and Training Strategy
To evaluate the benefit of multiband observations, two network configurations were implemented. The first is a single branch network processing input waveforms from either DECIGO or ET individually. The second is a dual branch fusion network designed for multiple detector inputs, where each branch independently processes data from one detector using identical ResNet substructures. Deep feature maps from both branches are concatenated in feature space and fused via a fully connected layer to form a unified representation. This late fusion approach preserves detector specific feature learning while leveraging the complementarity between the low frequency (DECIGO) and mid to high frequency (ET) bands, thereby enhancing discrimination of both interference–diffraction patterns and geometric magnification effects (Fig. 6).
@@
VI Results and Analysis
This section evaluates the model’s classification performance under different detector setups and SNR conditions. The analysis is divided into two parts. First, using datasets that span the full SNR range of unlensed signals (0100100), we evaluate and compare the classification results of models trained on individual detectors (ET and DECIGO) with those trained on combined data from both detectors. This comparison helps to quantify the effect of using multiple detectors on overall performance. Second, we examine how classification accuracy varies with SNR within the multiple detector configuration, to better understand how signal strength influences detection sensitivity and the reliability of lens classification.
VI.1 Single and Multiple Detector Performance
Using a comprehensive dataset spanning the 0100100 SNR range, we evaluated the model’s classification performance across five classes: PN, UL signals, and three representative lensing models (SIS, CIS, and NFW). Three input configurations were tested: single detector ET, single detector DECIGO, and the multiple detectors (ET+DECIGO). The dataset was divided into training and testing subsets in a 9:19{:}1 ratio. Each class contained $4\times 10^{4}$ samples, corresponding to $3.6\times 10^{4}$ training and $4\times 10^{3}$ testing samples per class, giving a total of $1.8\times 10^{5}$ training and $2\times 10^{4}$ testing samples.
@@ VI.1 Single and Multiple Detector Performance
$\mathrm{Micro~average}(M)=\frac{1}{C}\sum_{i=1}^{C}M_{i},$ (43)
@@
Comprehensive comparisons under different input configurations are summarised in Tables 35. The accuracy, average $F_{1}$, and average AUC trends are shown in Fig. 7, and the corresponding ROC curves and confusion matrices are presented in Fig. 8 and Fig. 9, respectively.
@@
The confusion matrix (Fig. 9(c)) exhibits strong diagonal dominance, with mutual misclassification between SIS and CIS below $2\%$, and NFW misclassification below $1\%$. The ROC curves (Fig. 8) show that the joint model completely encompasses both single detector models, with a micro average AUC approaching unity and individual class AUCs improving by about 0.100.100.250.25. This demonstrates a clear enhancement in both sensitivity and stability.
@@
Overall, based on the metrics in Tables 35 and the visual comparisons in Figs. 89, the ET+DECIGO joint model outperforms both single detector counterparts in accuracy, confusion structure, and robustness, confirming the physical and statistical advantages of multiband data fusion for detecting lensed GWs in noisy environments.
VI.2 Performance of the Multiple Detector at Different SNR
To evaluate the robustness and sensitivity of the multiple detector (ET+DECIGO) model under varying SNR, the model was trained and evaluated separately within five SNR intervals: 552020, 20204040, 40406060, 60608080, and 8080100100. Each interval included five classes (PN, UL, SIS, CIS, NFW) with a total of $2\times 10^{5}$ samples ($4\times 10^{4}$ per class). The data were divided into training and testing subsets with a 9:19{:}1 split. Model performance was quantified using Accuracy, Precision, Recall, $F_{1}$ score, and Macro Average metrics. Additionally, confusion matrices (Fig. 12) and ROC curves (Fig. 11) were plotted to visualise class separation across different SNR regimes. Quantitative metrics for each interval are summarised in Tables 610, and the evolution of accuracy, average $F_{1}$, and average AUC with increasing SNR is shown in Fig. 10.
@@ VI.2 Performance of the Multiple Detector at Different SNR
At the lowest SNR interval (552020), model performance is strongly degraded by noise contamination. The accuracy reaches only $0.6339$, and the macro averaged $F_{1}$ score is $0.6059$ (Table 6). While the “PN” class is perfectly identified ($F_{1}=1.000$), the “UL” and three lensed classes exhibit substantial confusion. The confusion matrix (Fig. 12(a)) shows that SIS and CIS suffer the most from mutual misclassification, with correct recognition rates of about $39\%$ and $15\%$, respectively, and a fraction of samples incorrectly labelled as NFW. This indicates that at low SNR, waveform interference features are heavily masked by noise, hindering the network’s ability to extract stable phase dependent structures. The ROC curves (Figs. 11(d), 11(e), and 11(f)) show that the AUC values for the three lensed classes remain in the range 0.840.840.850.85, suggesting partial robustness to noise but limited discriminative confidence. In this regime, the joint model’s advantage is not yet fully realised, being primarily constrained by the low SNR boundary of the input data.
@@
As the SNR increases to moderate levels (20204040 and 40406060), the model performance improves markedly. In the 20204040 range, the accuracy rises to $0.7074$ (Table 7), and the macro averaged $F_{1}$ score reaches $0.6427$. Both SIS and NFW exhibit substantial gains relative to the low SNR regime, with $F_{1,\text{SIS}}$ improving from $0.4431$ to $0.5891$ and $F_{1,\text{NFW}}$ from $0.5496$ to $0.6620$. The CIS class, however, remains poorly classified, with $F_{1}=0.0215$.
At SNRs of 40406060, performance increases sharply: accuracy reaches $0.9642$ (Table 8), and the macro averaged $F_{1}$ rises to $0.9634$. The five classes become nearly perfectly separated, with “PN” and “UL” achieving $F_{1}=1.000$. Among the lensed signals, $F_{1,\text{SIS}}=0.9170$, $F_{1,\text{CIS}}=0.9009$, and $F_{1,\text{NFW}}=0.9995$. The confusion matrix (Fig. 12(c)) exhibits clear diagonal dominance, with residual misclassifications below $20\%$. The ROC curves (Fig. 11) show AUC values exceeding $0.9$ for all classes, with an average AUC near $0.99$, confirming the model’s capacity to distinguish diffraction and geometric magnification features at moderate SNRs. The accuracy, average $F_{1}$, and average AUC in Fig. 10 displays a distinct transition around an SNR of $40$, indicating that multiple detector data fusion becomes particularly effective beyond this threshold.
@@
At high SNR (6060100100), model performance approaches saturation. In the 60608080 SNR interval, accuracy reaches $0.9983$ (Table 9), and the macro averaged $F_{1}$ is $0.9982$, with both Precision and Recall for all classes nearly unity. The confusion matrix (Fig. 12(d)) is almost perfectly diagonal, showing only negligible misclassifications.
In the 8080100100 SNR range, accuracy slightly decreases to $0.9903$ (Table 10), mainly due to minor boundary overlaps among classes, but remains close to ideal. The $F_{1}$ scores are $F_{\text{SIS}}=0.9750$, $F_{1,\text{CIS}}=0.9767$, and $F_{1,\text{NFW}}=0.9994$. The ROC curves (Figs. 11(a) and 11(b)) show $\mathrm{AUC}\approx 1.000$, confirming that the model achieves near perfect discrimination at high SNR. The confusion matrix (Fig. 12(e)) further verifies precise classification across all five classes, with virtually no off diagonal errors.
The overall trend (Fig. 10) reveals a rapid convergence of model performance with increasing SNR—from $63.4\%$ accuracy in the 552020 range to $99.8\%$ in the 60608080 interval—followed by stability at higher SNRs. These results confirm that the multiple detector configuration substantially enhances the extraction of interference and amplitude features at low to moderate SNRs, while maintaining calibrated probability outputs and precise classification at high SNR. Combining the results in Tables 610 and Figs. 1112, it is evident that the ET+DECIGO model achieves consistently high accuracy and robustness across all SNR regimes, highlighting the crucial role of multiband information fusion in detecting lensed GWs under noise conditions.
VII Conclusion and Outlook
@@
Future research can further extend and refine this framework in several directions. In data generation, incorporating more diverse lens models, such as ellipsoidal potentials [hawken2009gravitational] and double lens systems [collett2016compound, vegetti2024strong], may enhance the adaptability and generalization of the model in complex lensing environments. On the algorithmic side, future studies could investigate network architectures that integrate time and frequency domain representations, or introduce attention mechanisms to better capture long timescale features [chen2023long, brauwers2021general]. Joint task learning strategies may also enable simultaneous lens classification and parameter estimation, for example, predicting lens mass or impact parameter. The proposed approach can be further evaluated under more realistic observational conditions, including multiple overlapping sources, non stationary noise, and anisotropic detector responses, to examine its robustness and scalability. As third generation ground based detectors (ET) and space detectors (DECIGO, LISA [amaro2017laser], Taiji [ruan2020taiji], TianQin [luo2016tianqin]) continue to advance, multiband and multiband observations are expected to play an increasingly important role in gravitational wave astronomy. Within this context, the present framework offers a useful foundation for rapid identification of lensed GWs, inference of lens parameters, and statistical investigations related to dark matter and cosmology, contributing to the ongoing development of precision gravitational wave astrophysics.
@@ Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203004), the Fundamental Research Funds for the Central Universities Project (Grant No. 2024IAIS-ZD009), the National Natural Science Foundation of China (Grant Nos. 12575072 and 12347101), and the Natural Science Foundation of Chongqing (Grant No. CSTB2023NSCQ-MSX0103). This work made use of TensorFlow [Abadi_TensorFlow_Large-scale_machine_2015], PyCBC [alex_nitz_2024_10473621], GLoW_public [GLoW_public], and other open source Python packages.